# NestedSGX: Bootstrapping Trust to Enclaves within Confidential VMs

### v1

@@ 6.3. Real-world evaluations
Hash join. Hash join is used to implement the “equivalent-join” operation in modern databases. It involves creating a hash table from rows of the first table and then probes it with rows of the second table. We used the open-source implementation of the algorithm in SGXGauge (Kumar et al., 2022; sgx, 2023b). We varied the size of the first table (from 500K to 1M records) and fixed the second table (100K), effectively varying the memory and compute intensive nature of the workload. As shown in Figure. 6(a), the overhead of NestedSGX is negligible. In several cases, NestedSGX outperformed the native VM, likely due to the fact that the enclave’s gPAs were pre-allocated in NestedSGX, which avoided page faults.
SQLite. We utilized the publicly available version of SQLite (v.3.19.3) operating on Intel SGX (sql, 2021). We configured the database to function in-memory mode and incorporated the client within the enclave. We conducted performance evaluations across various record sizes, measuring the time required for 100,000 database operations. Specifically, we employed three representative workloads within the Yahoo! Cloud Serving Benchmark (YCSB) suite (ycs, 2021): workload $A$ (update heavy with 50% reads and 50% updates), $D$ (read latest, i.e., delete old ones, insert new ones and read mostly the new ones), and $F$ (short range scan) for the evaluation. As shown in Figure 6(b), NestedSGX introduces about 23.0% overhead on average over running in the native VM, and about 7.0% overhead on average over running in the simulation mode.
TLS server. We utilized the open-source SGX-OpenSSL project and a sample implementation of a TLS server in an enclave (sgx, 2023a) for our evaluation. Additionally, we employed OpenSSL to run a TLS server in the native VM for comparison purposes. The TLS client operated within the native CVM and sent requests to the server over the local loop-back. To measure the average latency, we employed the tls-perf benchmarking tool (tls, 2023). Figure 6(c) illustrates the results, indicating that NestedSGX incurs approximately a 35.9% overhead over running in the native VM and a 11.4% overhead over running in the simulation mode.
@@ 7. Discussions
Supporting other VM-based TEEs. The VMPL feature is not available on other TEEs, such as Intel TDX and ARM CCA. Essentially, VMPL partitions the CVM into separated VMs, each working on a part of gPA with different access permissions. NestedSGX can be transplanted to other TEEs, utilizing one CVM for the guest OS and App, and an additional one for the security monitor and enclaves. The CVMs are isolated from each other, similar to the VMPL isolation. The switches between them involve suspending the vCPU in one CVM and invokes the vCPU in the other CVM.
In SEV-SNP, memory sharing is straightforward. For instance, when emulating EADD, the security monitor can directly duplicate the page from the guest OS to the enclave and establish the enclave’s page table. In TDX and CCA, memory sharing can be achieved with the assistance of the TDX module and Realm Management Monitor (RMM). However, it is worth mentioning that the TDX module and RMM are critical for memory isolation and secure management of CVMs, and the customization of these components may not be allowed by chip companies. In fact, only the TDX module signed by Intel can run in the Secure Arbitration Mode (SEAM).
@@ 8. Related Works
However, it is important to note that IMA in general assumes the trust of the boot components, which includes the BIOS, bootloader and guest OS, leading to potentially large TCB and expanded attack surface. For example, it has been shown that IMA can be bypassed with a malicious block device (Bohling et al., 2020). In contrast, the core root of trust for measurement of NestedSGX (i.e., the security monitor) operates at VMPL0 and we do not place trust on the guest OS. Moreover, the security monitor not only measures the integrity of both enclave code and data but also ensures that the enclave runs in an isolated environment from the guest OS, which is beyond the design goal of IMA.
Establishing the enclave abstraction within TEEs. With the emergence of Intel SGX, endeavors have been undertaken to enhance the interoperability between Intel SGX and other TEE platforms. Notably, Komodo (Ferraiuolo et al., 2017), Sanctuary (Brasser et al., 2019), and SecTEE (Zhao et al., 2019) offer support for enclave abstractions on ARM TrustZone. vSGX (Zhao et al., 2022) effectively virtualizes Intel SGX on AMD SEV, ensuring complete binary compatibility and enabling the execution of unmodified SGX programs. HyperEnclave (Jia et al., 2022) re-introduces Intel SGX on standard AMD hardware through the use of a small and trusted hypervisor. In contrast to these works, NestedSGX sets itself apart by adopting a distinctive model. While the guest OS remains shielded from the untrusted hypervisor, it is not entirely trusted. Consequently, NestedSGX enables enhanced protection for enclaves within the feature-rich guest OS.
Security and performance enhancement using the VMPL feature. Hecate (Ge et al., 2022) utilizes VMPL0 as an L1 hypervisor, operating within the CVM on top of the untrusted L0-hypervisor (host VMM). Hecate enables the migration of on-premises workloads by maintaining compatibility with unmodified VM images and implementing security policies, such as network firewalls. Honeycomb (Mai et al., 2023) executes a validator within VMPL0, safeguarded from other software. The validator analyzes the binary code of a GPU kernel to validate that each memory instruction within the GPU kernel can only access specific regions of the virtual address space, employing static analysis techniques. SVSM-vTPM (Narayanan et al., 2023) employs VMPL to isolate the virtual TPM (vTPM) from the guest OS, therefore ensuring the integrity of vTPM’s functionalities and the security of the root key for remote attestation. However, all these solutions do not provide isolation for user programs from the untrusted guest OS.
@@ 9. Conclusions
Although confidential computing aims to protect the confidentiality and integrity of code and data within TEEs, it remains challenging to control code within CVMs since the feature-rich guest OS is free to load any code. In this paper, we present NestedSGX, a secure and efficient solution for maintaining control over the integrity of code and data in TEEs. By leveraging VMPL-based isolation, NestedSGX creates hardware enclaves inside guest VMs, which ensures remote attestation of trusted and measured code. NestedSGX is compatible with Intel SGX and easily integrated into existing systems. Our evaluation shows that NestedSGX incurs small overhead in most real-world applications.

### v2

@@ 6.3. Real-world evaluations
We conducted evaluations on 4 real-world applications, including hash join (memory- and computation-intensive tasks), SQLite (I/O- and memory-intensive tasks) and TLS server (I/O-intensive tasks) which ran directly on NestedSGX, and Redis (I/O- and memory-intensive tasks) which ran with Occlum on NestedSGX.
Hash join. Hash join is used to implement the “equivalent-join” operation in modern databases. It involves creating a hash table from rows of the first table and then probes it with rows of the second table. We used the open-source implementation of the algorithm in SGXGauge (Kumar et al., 2022; sgx, 2023b). We varied the size of the first table (from 500K to 1M records) and fixed the second table (100K), effectively varying the memory and computation intensive nature of the workload. As shown in Figure. 7(a), the overhead of NestedSGX is negligible.
SQLite. We utilized the publicly available version of SQLite (v3.19.3) operating on Intel SGX (sql, 2021). To evaluate the memory performance of NestedSGX, we configured the database to function in-memory mode and incorporated the client within the enclave. We conducted performance evaluations across various record sizes, measuring the time required for 100,000 database operations. Specifically, we employed three representative workloads within the Yahoo! Cloud Serving Benchmark (YCSB) suite (ycs, 2021): workload $A$ (update heavy with 50% reads and 50% updates), $D$ (read latest, i.e., delete old ones, insert new ones and read mostly the new ones), and $F$ (short range scan) for the evaluation. As shown in Figure 7(b), NestedSGX introduces about 0.77% overhead on average over the baseline.
TLS server. We utilized the open-source SGX-OpenSSL project and a sample implementation of a TLS server in an enclave (sgx, 2023a) for our evaluation. The TLS client operated within the native CVM and sent requests to the server over the local loop-back. To measure the average latency, we employed the tls-perf benchmarking tool (tls, 2023) and set the number of requesting threads to 1, 2, 5, 10, 20, 30, 40, 50 respectively. As illustrated in Figure 7(c), NestedSGX incurs approximately a 9.75% overhead over the baseline.
Redis. We ran a Redis (v6.0.9) database server with Occlum on NestedSGX. The database was configured as an in-memory setup, and we applied the YCSB workload A for our evaluation. We first loaded 5,000 records (in total 5 MB data), and subsequently executed 10,000 operations from 20 clients over the local loopback. We gradually increased the request frequency and measured the corresponding latency at different throughput levels. As depicted in Figure 7(d), NestedSGX achieved an approximate throughput of 84.32% (15.68% overhead) compared to the baseline. The latency was found to be 6.4% under low load conditions and increased to approximately 55% as the system approached its maximum throughput.
@@ 7. Discussions
Supporting other VM-based TEEs. Recently, ARM CCA introduces the support of different planes within a CVM, in which each plane is essentially a separated VM but they all share the same guest physical address space. Plane 0 is more privileged and can run a paravisor which manages the switches between other planes and can restrict the memory accessible by other planes (Arm, 2024). According to ARM’s roadmap, the paravisor is designed to support secure services such as vTPM emulation. Similarly, we can re-purpose the paravisor to support running userspace enclaves. Similarly, TDX supports memory sharing among CVMs with the assistance of the TDX module. Therefore, it is promising that NestedSGX can be extended to other CVM platforms. However, it is worth mentioning that the TDX module and CCA’s Realm Management Monitor (RMM) are critical for the memory isolation and secure management of CVMs, and the customization of these components may not be allowed by chip companies. In fact, only the TDX module signed by Intel can run in the Secure Arbitration Mode (SEAM).
@@ 8. Related Works
However, IMA in general assumes the trust of the boot components, which includes the BIOS, bootloader and guest OS, leading to potentially large TCB and expanded attack surface. For example, it has been shown that IMA can be bypassed with a malicious block device (Bohling et al., 2020). In contrast, the core root of trust for measurement of NestedSGX (i.e., the security monitor) operates at VMPL0 and we do not place trust on the guest OS. Moreover, the security monitor not only measures the integrity of both enclave code and data but also ensures that the enclave runs in an isolated environment from the guest OS, which is beyond the design goal of IMA.
Establishing the enclave abstraction within TEEs. With the emergence of Intel SGX, endeavors have been undertaken to enhance the interoperability between Intel SGX and other TEE platforms. Notably, Komodo (Ferraiuolo et al., 2017), Sanctuary (Brasser et al., 2019), and SecTEE (Zhao et al., 2019) offer support for enclave abstractions on ARM TrustZone. HyperEnclave (Jia et al., 2022) re-introduces Intel SGX on standard AMD hardware through the use of a small and trusted hypervisor. In contrast to these works, NestedSGX sets itself apart by adopting a distinctive model. While the guest OS remains shielded from the untrusted hypervisor, it is not entirely trusted. Consequently, NestedSGX enables enhanced protection for enclaves within the feature-rich guest OS. Most related to our work are vSGX (Zhao et al., 2022) and Veil (Ahmad et al., 2024). A more detailed comparison with them is provided in Sec. 3.1.
Security and performance enhancement using the VMPL feature. Hecate (Ge et al., 2022) utilizes VMPL0 as an L1 hypervisor, operating within the CVM on top of the untrusted L0-hypervisor (host VMM). Hecate enables the migration of on-premises workloads by maintaining compatibility with unmodified VM images and implementing security policies, such as network firewalls. Honeycomb (Mai et al., 2023) executes a validator within VMPL0, safeguarded from other software. The validator analyzes the binary code of a GPU kernel to validate that each memory instruction within the GPU kernel can only access specific regions of the virtual address space, employing static analysis techniques. SVSM-vTPM (Narayanan et al., 2023) employs VMPL to isolate the virtual TPM (vTPM) from the guest OS, ensuring the integrity of vTPM’s functionalities and the security of the root key for remote attestation. However, all these solutions do not provide isolation for user programs from the untrusted guest OS.
@@ 9. Conclusions
Although confidential computing aims to protect the confidentiality and integrity of code and data within TEEs, it remains challenging to control code within CVMs since the feature-rich guest OS is free to load any code. This paper presents NestedSGX, a secure and efficient solution for maintaining control over the integrity of code and data in TEEs. NestedSGX creates hardware enclaves inside guest VMs with VMPL, which ensures remote attestation of trusted and measured code. NestedSGX is compatible with Intel SGX and easily integrated into existing systems. Our evaluation shows NestedSGX incurs small overhead in most real-world applications.